OpenStudy (anonymous):

There are 6 blue marbles,7 black marbles,4 orange marbles,and 3 green marbles in a bag. Once a marble is drawn,it is REPLACED. Find the probability of the outcome heres the question a blue marble,an orange marble,and then a green marble a)0 b)9/1000 c)6/35 d)3/500

OpenStudy (anonymous):

Blue is 6 out of the 20 marbles Orange is 4 out of the 20 marbles Green is 3 out of the 20 marbles Probability is 6/20 * 4/20 * 3/20 I would reduce the 6/20 to 3/10 and the 4/20 to 1/5 so Probability = 3/10 * 1/5 * 3/20 = 9/1000 So you would get a blue then an orange then a green approximately 9 out of every 1000 tries.

OpenStudy (anonymous):

Thanksss soooo much im so tired of trying to do this and not remembering im happy my teacher let me take this home to do i go 9 worong outa 20

OpenStudy (anonymous):

You r welcome

OpenStudy (anonymous):

You only do it this way when you "replace" after each draw.

OpenStudy (anonymous):

lol im in 7th grade and i have AIMS the week after this ugh